# Fano threefolds in positive characteristic III

### v3

@@ 5.3. Case $C-D$
By Lemma 4.15(1), $f$ is a finite morphism. It follows from $L_{1}\simeq\mathcal{O}_{\mathbb{P}^{2}}(1)$ and $L_{2}\simeq\mathcal{O}_{\mathbb{P}^{1}}(1)$ that $\deg f=(f_{1}^{*}L_{1})^{2}\cdot f_{2}^{*}L_{2}$. By Corollary 4.5, any fibre of $f_{1}:X\to Y_{1}=\mathbb{P}^{2}$ is numerically equivalent to $\frac{2}{\mu_{1}}\ell_{1}$, i.e., $(f_{1}^{*}L_{1})^{2}\equiv\frac{2}{\mu_{1}}\ell_{1}$. We then obtain
@@
(2) Assume that $R_{2}$ is of type $C_{2}$. Then we have $\mu_{1}=1$ and $\deg f=1$. Hence, $f:X\to\mathbb{P}^{2}\times\mathbb{P}^{1}$ is a finite birational morphism of normal varieties, which is automatically an isomorphism. In particular, $R_{2}$ is of type $D_{3}$. ∎
@@
Let $X$ be a Fano threefold with $\rho(X)=2$. Let $R_{1}$ and $R_{2}$ be extremal rays of ${\operatorname{NE}}(X)$. Assume that $R_{1}$ is of type $C$ and $R_{2}$ is of type $D$. Then one and only one of the following holds.
@@
    $R_{1}$ is of type $C_{1}$, $R_{2}$ is of type $D_{1}$, $(-K_{X})^{3}=6$, and there exists a split double cover $f:X\to\mathbb{P}^{2}\times\mathbb{P}^{1}$ such that $(f_{*}\mathcal{O}_{X}/\mathcal{O}_{\mathbb{P}^{2}\times\mathbb{P}^{1}})^{-1}\simeq\mathcal{O}_{\mathbb{P}^{2}\times\mathbb{P}^{1}}(2,1)$.
@@
    $R_{1}$ is of type $C_{1}$, $R_{2}$ is of type $D_{2}$, $(-K_{X})^{3}=24$, and there exists a split double cover $f:X\to\mathbb{P}^{2}\times\mathbb{P}^{1}$ such that $(f_{*}\mathcal{O}_{X}/\mathcal{O}_{\mathbb{P}^{2}\times\mathbb{P}^{1}})^{-1}\simeq\mathcal{O}_{\mathbb{P}^{2}\times\mathbb{P}^{1}}(1,1)$.
@@
    $R_{1}$ is of type $C_{2}$, $R_{2}$ is of type $D_{3}$, $(-K_{X})^{3}=54$, and $X\simeq\mathbb{P}^{2}\times\mathbb{P}^{1}$.
@@ 5.4. Case $C-E$ (primitive)
Notation 5.24.
We use Notation 5.1. Assume that $R_{1}$ is of type $C$ and $R_{2}$ is of type $E_{2},E_{3},E_{4}$, or $E_{5}$. Set $D:={\operatorname{Ex}}(f_{2}:X\to Y_{2})$, which is a prime divisor on $X$.
Let us consider the induced morphism $f_{1}|_{D}\colon D\to Y_{1}=\mathbb{P}^{2}$.
@@
We use Notation 5.24. Then the following hold.
@@
Since $R_{2}$ is of type $E_{2},E_{3},E_{4}$ or $E_{5}$, $f_{2}(D)$ is a point on $Y$. Therefore, we get
@@ 5.4.1. Case $C_{1}-E$
We use Notation 5.26. Then $R^{1}(f_{1})_{*}\mathcal{O}_{X}(-K_{X}-D)=0$ and there exists an exact sequence:
@@
We use Notation 5.26. Then $(f_{1}|_{D})^{*}\mathcal{O}_{\mathbb{P}^{2}}(n)\simeq\mathcal{O}_{D}(n)$ for all $n\in\mathbb{Z}$.
@@
We use Notation 5.26. Then the following hold.
@@
Let us show (1). For a closed point $t\in\mathbb{P}^{2}$ and the fibre $X_{t}$ of $f_{1}:X\to Y_{1}=\mathbb{P}^{2}$ over $t$, we have $-K_{X}\cdot X_{t}=D\cdot X_{t}=2$. Hence $h^{0}(X_{t},\mathcal{O}_{X_{t}}(-K_{X}-D))=1$ for all $t\in Y_{1}=\mathbb{P}^{2}$. By Grauert’s theorem [Har77, Ch. III, Theorem 12.9], $(f_{1})_{*}(\mathcal{O}_{X}(-K_{X}-D))$ is an invertible sheaf on $\mathbb{P}^{2}$. Hence we can write
@@
We use Notation 5.26. Then $\mathcal{O}_{P}(X)\simeq\mathcal{O}_{P}(2)$, where $\mathcal{O}_{P}(1)$ denotes the tautological line bundle of the $\mathbb{P}^{2}$-bundle structure $\pi:P=\mathbb{P}((f_{1})_{*}\mathcal{O}_{X}(-K_{X}))\to\mathbb{P}^{2}$ and we set $\mathcal{O}_{P}(2):=\mathcal{O}_{P}(1)^{\otimes 2}$.
@@
Recall that we have the following diagram (Notation 5.26, Proposition 5.29):
@@
Taking the intersection number of this equation with a general fibre of $\pi|_{X}\colon X\to\mathbb{P}^{2}$, we have $1=3-n$, i.e., $n=2$. Hence we obtain $\mathcal{O}_{X}\simeq\pi^{*}\mathcal{O}_{\mathbb{P}^{2}}(a)|_{X}$. We see that $a=0$ by taking the intersection number of this equation with a curve $C$ on $X$ such that $\pi(C)$ is a curve. ∎
@@
It follows from Proposition 7.1(5) that
@@
We use Notation 5.26. Then there exist a split double cover
@@
such that for any closed point $z\in\mathbb{P}^{2}=Z$, the base change of (5.32.1) by $z\hookrightarrow\mathbb{P}^{2}=Z$ can be written as follows
@@
Suppose $X\cap S\neq\emptyset$. Note that $P=\mathbb{P}_{\mathbb{P}^{2}}(\mathcal{O}_{\mathbb{P}^{2}}\oplus\mathcal{O}_{\mathbb{P}^{2}}(1)\oplus\mathcal{O}_{\mathbb{P}^{2}}(2))$ is covered by $\mathbb{A}^{4}$. Then it follows from [Har77, Ch. I, Proposition 7.1] that $\dim(X\cap S)\geq 3+2-4=1$. Hence there exists a curve $C$ with $C\subset X\cap S$. By $\mathcal{O}_{X}(-K_{X})\simeq\mathcal{O}_{P}(1)|_{X}$ and $\mathcal{O}_{P}(1)|_{S}\simeq\mathcal{O}_{S}$, we obtain
@@ 5.4.2. Case $C_{2}-E$
Let $X$ be a Fano threefold with $\rho(X)=2$. Let $R_{1}$ and $R_{2}$ be extremal rays of ${\operatorname{NE}}(X)$. Assume that $R_{1}$ is of type $C$ and $R_{2}$ is of type $E_{2},E_{3},E_{4}$, or $E_{5}$. Then one and only one of the following holds.
@@
    $R_{1}$ is of type $C_{1}$, $R_{2}$ is of type $E_{3}$ or $E_{4}$, $(-K_{X})^{3}=14$, and there exists a split double cover $f:X\to V_{7}$ such that $f_{*}\mathcal{O}_{X}/\mathcal{O}_{V_{7}}\simeq\mathcal{O}_{V_{7}}(-\frac{1}{2}K_{V_{7}})^{-1}$.
@@
    $R_{1}$ is of type $C_{2}$, $R_{2}$ is of type $E_{2}$, $(-K_{X})^{3}=56$, and $X\simeq\mathbb{P}(\mathcal{O}_{\mathbb{P}^{2}}\oplus\mathcal{O}_{\mathbb{P}^{2}}(1))$.
@@
    $R_{1}$ is of type $C_{2}$, $R_{2}$ is of type $E_{5}$, $(-K_{X})^{3}=62$, and $X\simeq\mathbb{P}(\mathcal{O}_{\mathbb{P}^{2}}\oplus\mathcal{O}_{\mathbb{P}^{2}}(2))$.
@@
The assertion follows from Lemma 5.25, Proposition 5.32, and Lemma 5.33. ∎
@@ 5.5. Case $E_{1}-C$
    $(-K_{X})^{3}=18,r_{1}=2,(-K_{X_{1}})^{3}=24,\deg B_{1}=1,g(B_{1})=0,\deg\Delta_{f_{2}}=5$ (No. 2-11).
@@
    $(-K_{X})^{3}=22,r_{1}=2,(-K_{X_{1}})^{3}=32,\deg B_{1}=2,g(B_{1})=0,\deg\Delta_{f_{2}}=4$ (No. 2-16).
@@
    $(-K_{X})^{3}=26,r_{1}=2,(-K_{X_{1}})^{3}=40,\deg B_{1}=3,g(B_{1})=0,\deg\Delta_{f_{2}}=3$ (No. 2-20).
@@ 5.6. Case $E_{1}-D$
$(-K_{X})^{3}=(3H_{1}+H_{2})^{3}=27H_{1}^{3}+27H_{1}^{2}\cdot H_{2}+9H_{1}\cdot H_{2}^{2}+H_{3}^{3}=27+27+0+0=54.$
@@
We have $d_{2}=8$. By Lemma 5.39, the following hold:
@@
By (B), $r_{1}\geq 3$. Moreover, we get
@@
It holds that
@@
Take two fibres $F_{1}$ and $F_{2}$ of $\pi:X\to\mathbb{P}^{1}$. Set $D_{1}:=\sigma_{*}F_{1}$ and $D_{2}:=\sigma_{*}F_{2}$. For each $i\in\{1,2\}$, we get $D_{i}=\sigma_{*}F_{i}\sim\sigma_{*}F\overset{{\rm(ii)}}{\sim}D$. We then obtain $\Gamma\supset(D_{1}\cap D_{2})_{{\operatorname{red}}}$. Since $D_{1}\cap D_{2}$ is of pure one-dimensional, we get $(D_{1}\cap D_{2})_{{\operatorname{red}}}=\Gamma$ or $(D_{1}\cap D_{2})_{{\operatorname{red}}}=\emptyset$. We then obtain $(D_{1}\cap D_{2})_{{\operatorname{red}}}=\Gamma$, as otherwise the equality $(D_{1}\cap D_{2})_{{\operatorname{red}}}=\emptyset$ would lead to the following contradiction:
@@
It is enough to show the scheme-theoretic equality $D_{1}\cap D_{2}=\Gamma$. Recall that we have the equality $(D_{1}\cap D_{2})_{{\operatorname{red}}}=\Gamma$. Since $D_{1}\cap D_{2}$ is Cohen-Macaulay, it suffices to prove that $D_{1}\cap D_{2}$ is $R_{0}$, i.e., $\mathcal{O}_{D_{1}\cap D_{2},\xi}$ is reduced for the generic point $\xi$ of $D_{1}\cap D_{2}$. This follows from $(-K_{Y})\cdot(D_{1}\cap D_{2})=-K_{Y}\cdot\Gamma$, because the effective $1$-cycle associated with $D_{1}\cap D_{2}$ is determined by its length at the generic point [Bad01, Lemma 1.18]. ∎

### v4

@@ 5.3. Case $C-D$
By Lemma 4.14(1), $f$ is a finite morphism. It follows from $L_{1}\simeq\mathcal{O}_{\mathbb{P}^{2}}(1)$ and $L_{2}\simeq\mathcal{O}_{\mathbb{P}^{1}}(1)$ that $\deg f=(f_{1}^{*}L_{1})^{2}\cdot f_{2}^{*}L_{2}$. By Corollary 4.5, every fibre of $f_{1}:X\to Y_{1}=\mathbb{P}^{2}$ is numerically equivalent to $\frac{2}{\mu_{1}}\ell_{1}$, i.e., $(f_{1}^{*}L_{1})^{2}\equiv\frac{2}{\mu_{1}}\ell_{1}$. We then obtain
@@
(2) Assume that ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}R_{1}}$ is of type $C_{2}$. Then we have $\mu_{1}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}$ and $\deg f=1$. Hence, $f:X\to\mathbb{P}^{2}\times\mathbb{P}^{1}$ is a finite birational morphism of normal varieties, which is automatically an isomorphism. In particular, $R_{2}$ is of type $D_{3}$. ∎
@@
Let $X$ be a Fano threefold with $\rho(X)=2$. Let $R_{1}$ and $R_{2}$ be the extremal rays of ${\operatorname{NE}}(X)$. Assume that $R_{1}$ is of type $C$ and $R_{2}$ is of type $D$. Then one and only one of the following holds.
@@
    $R_{1}$ is of type $C_{1}$, $R_{2}$ is of type $D_{1}$, $(-K_{X})^{3}=6$, and there exists a split double cover $f:X\to\mathbb{P}^{2}\times\mathbb{P}^{1}$ such that $(f_{*}\mathcal{O}_{X}/\mathcal{O}_{\mathbb{P}^{2}\times\mathbb{P}^{1}})^{-1}\simeq\mathcal{O}_{\mathbb{P}^{2}\times\mathbb{P}^{1}}(2,1)$ (No. 2-2).
@@
    $R_{1}$ is of type $C_{1}$, $R_{2}$ is of type $D_{2}$, $(-K_{X})^{3}=24$, and there exists a split double cover $f:X\to\mathbb{P}^{2}\times\mathbb{P}^{1}$ such that $(f_{*}\mathcal{O}_{X}/\mathcal{O}_{\mathbb{P}^{2}\times\mathbb{P}^{1}})^{-1}\simeq\mathcal{O}_{\mathbb{P}^{2}\times\mathbb{P}^{1}}(1,1)$ (No. 2-18).
@@
    $R_{1}$ is of type $C_{2}$, $R_{2}$ is of type $D_{3}$, $(-K_{X})^{3}=54$, and $X\simeq\mathbb{P}^{2}\times\mathbb{P}^{1}$ (No. 2-34).
@@ 5.4. Case $C-E$ (primitive)
We use Notation 5.1. Assume that $R_{1}$ is of type $C$ and $R_{2}$ is of type $E_{2},E_{3},E_{4}$, or $E_{5}$. Set $D:={\operatorname{Ex}}(f_{2}:X\to Y_{2})$, which is a prime divisor on $X$. Then the following hold.
@@
Since $R_{2}$ is of type $E_{2},E_{3},E_{4}$ or $E_{5}$, $f_{2}(D)$ is a point on ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Y_{2}}$. Therefore, we get
@@ 5.4.1. Case $C_{1}-E$ (primitive)
We use Notation 5.25. Then $R^{1}(f_{1})_{*}\mathcal{O}_{X}(-K_{X}-D)=0$ and there exists an exact sequence:
@@
We use Notation 5.25. Then $(f_{1}|_{D})^{*}\mathcal{O}_{\mathbb{P}^{2}}(n)\simeq\mathcal{O}_{D}(n)$ for all $n\in\mathbb{Z}$.
@@
We use Notation 5.25. Then the following hold.
@@
Let us show (1). For a closed point $t\in\mathbb{P}^{2}$ and the fibre $X_{t}$ of $f_{1}:X\to Y_{1}=\mathbb{P}^{2}$ over $t$, we have $-K_{X}\cdot X_{t}=D\cdot X_{t}=2$. Hence $h^{0}(X_{t},\mathcal{O}_{X_{t}}(-K_{X}-D))=1$ for all $t\in Y_{1}=\mathbb{P}^{2}$. By Grauert’s theorem [Har77, Ch. III, Theorem 12.9], $(f_{1})_{*}\mathcal{O}_{X}(-K_{X}-D)$ is an invertible sheaf on $\mathbb{P}^{2}$. Hence we can write
@@
We use Notation 5.25. Then $\mathcal{O}_{X}(-K_{X})\simeq\mathcal{O}_{P}(1)|_{X}$ and $\mathcal{O}_{P}(X)\simeq\mathcal{O}_{P}(2)$, where $\mathcal{O}_{P}(1)$ denotes the tautological line bundle of the $\mathbb{P}^{2}$-bundle structure $\pi:P=\mathbb{P}((f_{1})_{*}\mathcal{O}_{X}(-K_{X}))\to\mathbb{P}^{2}$ and we set $\mathcal{O}_{P}(2):=\mathcal{O}_{P}(1)^{\otimes 2}$.
@@
Recall that we have the following diagram (Notation 5.25, Proposition 5.28):
@@
Taking the intersection number of this equation with a general fibre of $\pi|_{X}\colon X\to\mathbb{P}^{2}$, we have $1=3-n$, i.e., $n=2$. Hence we obtain $\mathcal{O}_{X}\simeq\pi^{*}\mathcal{O}_{\mathbb{P}^{2}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}a)|_{X}$. We see that $a=0$ by taking the intersection number of this equation with a curve $C$ on $X$ such that $\pi(C)$ is a curve. ∎
@@
Set $\mathcal{E}:=\mathcal{O}_{\mathbb{P}^{2}}\oplus\mathcal{O}_{\mathbb{P}^{2}}(1)\oplus\mathcal{O}_{\mathbb{P}^{2}}(2)$. It follows from Proposition 7.1(5) that
@@
We use Notation 5.25. Then there exist a split double cover
@@
such that for every closed point $z\in\mathbb{P}^{2}=Z$, the base change of (5.31.1) by $z\hookrightarrow\mathbb{P}^{2}=Z$ can be written as follows
@@
Suppose $X\cap S\neq\emptyset$. Since $P=\mathbb{P}_{\mathbb{P}^{2}}(\mathcal{O}_{\mathbb{P}^{2}}\oplus\mathcal{O}_{\mathbb{P}^{2}}(1)\oplus\mathcal{O}_{\mathbb{P}^{2}}(2))$ is smooth, we have ${\rm codim}_{P}\,T\leq{\rm codim}_{P}\,X+{\rm codim}_{P}\,S=1+2=3$ for an irreducible component $T$ of $X\cap S$ [EH16, Theorem 0.2]. Hence there exists a curve $C$ with $C\subset X\cap S$. By $\mathcal{O}_{X}(-K_{X})\simeq\mathcal{O}_{P}(1)|_{X}$ (Lemma 5.29) and $\mathcal{O}_{P}(1)|_{S}\simeq\mathcal{O}_{S}$, we obtain
@@ 5.4.2. Case $C_{2}-E$ (primitive)
Let $X$ be a Fano threefold with $\rho(X)=2$. Let $R_{1}$ and $R_{2}$ be the extremal rays of ${\operatorname{NE}}(X)$. Assume that $R_{1}$ is of type $C$ and $R_{2}$ is of type $E_{2},E_{3},E_{4}$, or $E_{5}$. Then one and only one of the following holds.
@@
    $R_{1}$ is of type $C_{1}$, $R_{2}$ is of type $E_{3}$ or $E_{4}$, $(-K_{X})^{3}=14$, and there exists a split double cover $f:X\to V_{7}$ such that $f_{*}\mathcal{O}_{X}/\mathcal{O}_{V_{7}}\simeq\mathcal{O}_{V_{7}}(-\frac{1}{2}K_{V_{7}})^{-1}$ (No. 2-8).
@@
    $R_{1}$ is of type $C_{2}$, $R_{2}$ is of type $E_{2}$, $(-K_{X})^{3}=56$, and $X\simeq\mathbb{P}(\mathcal{O}_{\mathbb{P}^{2}}\oplus\mathcal{O}_{\mathbb{P}^{2}}(1))$ (No. 2-35).
@@
    $R_{1}$ is of type $C_{2}$, $R_{2}$ is of type $E_{5}$, $(-K_{X})^{3}=62$, and $X\simeq\mathbb{P}(\mathcal{O}_{\mathbb{P}^{2}}\oplus\mathcal{O}_{\mathbb{P}^{2}}(2))$ (No. 2-36).
@@
The assertion follows from Lemma 5.24, Lemma 5.30, Proposition 5.31, and Lemma 5.32. ∎
@@ 5.5. Case $E_{1}-C$
    $(-K_{X})^{3}=18,r_{1}=2,(-K_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Y_{1}}})^{3}=24,\deg B_{1}=1,g(B_{1})=0,\deg\Delta_{f_{2}}=5$ (No. 2-11).
@@
    $(-K_{X})^{3}=22,r_{1}=2,(-K_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Y_{1}}})^{3}=32,\deg B_{1}=2,g(B_{1})=0,\deg\Delta_{f_{2}}=4$ (No. 2-16).
@@
    $(-K_{X})^{3}=26,r_{1}=2,(-K_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Y_{1}}})^{3}=40,\deg B_{1}=3,g(B_{1})=0,\deg\Delta_{f_{2}}=3$ (No. 2-20).
@@ 5.6. Case $E_{1}-D$
$(-K_{X})^{3}=(3H_{1}+H_{2})^{3}=27H_{1}^{3}+27H_{1}^{2}\cdot H_{2}+9H_{1}\cdot H_{2}^{2}+H_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}}^{3}=27+27+0+0=54.$
@@
We have $d_{2}=(-K_{X})^{2}\cdot H_{2}=8$. By Lemma 5.38, the following hold:
@@
By (B), ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r_{1}\in\{3,4\}}$. Moreover, we get
@@
By (C) (or Proposition 5.8), we obtain $r_{1}\in\{2,3,4\}$. It holds that
@@
Take two fibres $F_{1}$ and $F_{2}$ of $\pi:X\to\mathbb{P}^{1}$. Set $D_{1}:=\sigma_{*}F_{1}$ and $D_{2}:=\sigma_{*}F_{2}$. For each $i\in\{1,2\}$, we get $D_{i}=\sigma_{*}F_{i}\sim\sigma_{*}F\overset{{\rm(ii)}}{\sim}D$. We then obtain $\Gamma\supset(D_{1}\cap D_{2})_{{\operatorname{red}}}$. Since $D_{1}\cap D_{2}$ is of pure dimension one, we get $(D_{1}\cap D_{2})_{{\operatorname{red}}}=\Gamma$ or $(D_{1}\cap D_{2})_{{\operatorname{red}}}=\emptyset$. We then obtain $(D_{1}\cap D_{2})_{{\operatorname{red}}}=\Gamma$, as otherwise the equality $(D_{1}\cap D_{2})_{{\operatorname{red}}}=\emptyset$ would lead to the following contradiction:
@@
It is enough to show the scheme-theoretic equality $D_{1}\cap D_{2}=\Gamma$. Recall that we have the equality $(D_{1}\cap D_{2})_{{\operatorname{red}}}=\Gamma$. Since $D_{1}\cap D_{2}$ is Cohen-Macaulay, it suffices to prove that $D_{1}\cap D_{2}$ is $R_{0}$, i.e., $\mathcal{O}_{D_{1}\cap D_{2},\xi}$ is reduced for the generic point $\xi$ of $D_{1}\cap D_{2}$. This follows from $(-K_{Y})\cdot(D_{1}\cap D_{2})=(-K_{Y})\cdot D^{2}\overset{{\rm(i)}}{=}-K_{Y}\cdot\Gamma$, because the effective $1$-cycle associated with $D_{1}\cap D_{2}$ is determined by its length at the generic point [Bad01, Lemma 1.18]. ∎